what do you do to solve an equation when the solution is given as x is not equall to a number the sign was presented as an = sign with a / through it
thank you

Answer by funmath(2933)

You solve it the same way that you do when the = sign doesn't have a line through it. You're just finding out what x doesn't = now.
For example:
x+4=5
x+4-4=5-5
x=1
:
x+4(not=)5
x+4-4(not=)5-4
x(not=)1
:
Happy Calculating!!!